OpenStudy (anonymous):

Length: 2 and 1 over 4 feet Width: 1 foot Height: 1 and 1 over 4 feet The lengths of the sides of a small cube are 1 over 4 foot each. Part A: How many small cubes can be packed in the rectangular prism? Show your work. Part B: Use the answer obtained in part A to find the volume of the rectangular prism in terms of the small cube and a unit cube.

OpenStudy (anonymous):

@imqwerty

OpenStudy (anonymous):

@triciaal @Zeronknight

OpenStudy (mathmale):

Wouldn't it be easier to express "Length: 2 and 1 over 4 feet Width: 1 foot Height: 1 and 1 over 4 feet The lengths of the sides of a small cube are 1 over 4 foot each" as L=2 1/4 ft; W= 1 ft H= 1 1/4 ft small L = 1/4 ft ??

OpenStudy (anonymous):

either way i fine

OpenStudy (mathmale):

Note that this problem becomes easier if you convert these mixed numbers to improper fractions. H = 1 1/4 ft is equivalent to 5/4 ft.

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

So, what's your next step?

OpenStudy (anonymous):

ummmm my lil sis say idk

OpenStudy (mathmale):

1. Find the volume of the box of given dimensions. 2. Find the volume of a typical "small cube." 3. ?

OpenStudy (mathmale):

Volume of a rectangular box is L*W*H. You already know L, W and H. Multiply them together. Show all work, please.

OpenStudy (anonymous):

umm okay ill give it too her to show it

OpenStudy (anonymous):



OpenStudy (mathmale):

Caution: Convert L=2 1/4 ft; W= 1 ft H= 1 1/4 ft ...to improper fractions first.

OpenStudy (anonymous):

oh...

OpenStudy (mathmale):

Here's an example: 1 1/4 is the same as 1/4 + 1/4 + 1/4 + 1/4 + 1/4 = ? / 4

OpenStudy (mathmale):

What I'd do would be to multiply the first 1 in 1 1/4 by 4, add the 2nd 1 and write 5/4.

OpenStudy (mathmale):

that's the height, H.

OpenStudy (mathmale):

2 1/4: Same procedure. take the denom (4) and multiply the '2' by it. Result: 8. Can you finish this? 2 1/4 = ? / 4

OpenStudy (mathmale):

In effect, this gives us 4(2) 1 --- + --- = ?? 4 4

OpenStudy (anonymous):

um she say uhh uhhh uhhhh :O shes crying hold on

OpenStudy (mathmale):

Well, try this: L = 2 1/4 = 9/4 (ft) W = 1 (ft) H = 1 1/4 = 5/4 (ft) Volume of big box = L * W * H \[V=L*W*H=(\frac{ 9 }{ 4 })*(1)*(\frac{ 5 }{ 4 }) = ???\]

OpenStudy (mathmale):

the units of measurement here are (inches)^3, or \[inches^3\] .

OpenStudy (anonymous):

ohh i know she thanks you u help her a lot :) ill medal u

OpenStudy (mathmale):

sorry this process has been painful for your sister. Good luck!

OpenStudy (anonymous):

thanks